OpenStudy (anonymous):

Hey! Assuming that the emission line used for zinc determinations is 214 nm, and there are 4.50*10^10 atoms of zinc emitting light in an instrument flame at any instant, what energy (joules) must the flame continuously provide to achieve this level of emission? I know you know you want to answer it :p

OpenStudy (anonymous):

when you heat an atom sufficiently some of its electrons will jump to a higher energy called an excited state then the electrons will drop back to ground state releasing the same energy that caused the jump the energy is released as a photon atomic emission spectroscope is the instrument used to measure the wavelength and intensity of these emissions the energy of a single photon is found by E=h*f where h is Planck's constant and f is frequency but since c = f L, then f = c/L where c is speed of light, f is frequency and L is wavelength this makes the energy equation E = h*c/L so the energy of a single photon at 214 nm or 214E-9 m (we must use meters) is E = 3.00E8(6.626E-34)/214E-9 = 9.29E-19 J that's for just one photon so 9.29E-19 J/photon * 4.50E10 photons = 4.18E-8 J (assuming one electron jump per atom)

OpenStudy (anonymous):

i answered your question so could you answer mine? the internet isnt very helpful

OpenStudy (anonymous):

haha, I gave it a shot...I'm not too pro at Vectors so I tried a simple approach